I would like to know the intersection points of a parabola and line when:
y=x^2+x-3
y=x-1
Thank you.

Answer by Cromlix(4381)

Make the equations equal one another.
x^2 + x - 3 = x - 1
Combine the two equations
x^2 + x - x - 3 + 1 = 0
x^2 - 2 = 0
x^2 = 2
x = square root of two
Therefore the intersections with the
line is -square root 2 and + square root 2
Put these two values into line equation
y = x - 1
The y coords: -> (- square root 2 -1 = -2.41)
(+ square root 2 - 1 = 0.41)
(- square root 2, -2,41) (+ square root 2, 0.41)